# Sparse and Faithful Explanations Without Sparse Models

### 04 Mar 2024 v2

@@ Abstract
Even if a model is not globally sparse, it is possible for decisions made from that model to be accurately and faithfully described by a small number of features. For instance, an application for a large loan might be denied to someone because they have no credit history, which overwhelms any evidence towards their creditworthiness. In this work, we introduce the Sparse Explanation Value (SEV), a new way of measuring sparsity in machine learning models. In the loan denial example above, the SEV is 1 because only one factor is needed to explain why the loan was denied. SEV is a measure of decision sparsity rather than overall model sparsity, and we are able to show that many machine learning models – even if they are not sparse – actually have low decision sparsity, as measured by SEV. SEV is defined using movements over a hypercube, allowing SEV to be defined consistently over various model classes, with movement restrictions reflecting real-world constraints. We proposed the algorithms that reduce SEV without sacrificing accuracy, providing sparse and completely faithful explanations, even without globally sparse models.
@@ 1 Introduction
As machine learning is increasingly leveraged in business and societal contexts to make important decisions, there is an increasing need for insight into why these decisions were made. One classical important measure of interpretability is sparsity; famously, humans can handle only $7\pm 2$ cognitive entities at once (Miller, 1956). Traditionally, sparsity is a property of the entire model, which we call global sparsity, for instance, the total number of terms in linear models, the total number of parameters in neural networks, or the number of leaves in decision trees (Murdoch et al., 2019). In this work, we argue that this notion of sparsity, while undeniably useful, may be overly restrictive. This is because we do not need the model to be globally sparse in order to generate a sparse explanation for use in decision-making. A far more relevant desideratum for users is that the explanation for each individual prediction is sparse. By this, we mean that the local explanation for a specific prediction depends only on a small number of features, which may differ across units. Since individuals care only about decisions made about themselves, we should consider aiming for sparse explanations – and not necessarily sparse models. That is, having an explanation for each prediction that does not necessarily contain all factors used by the model. Having sparse explanations, like providing only 1-3 features for explanations for each predicted instance, allows for increased transparency in model-assisted decision-making processes, even if the model is not globally sparse.
@@
For example, consider a scenario where a loan applicant is denied a large loan due to their lack of credit history. This denial might occur even though the applicant has several positive factors, such as having a bank account, being employed, and not having a criminal record. In this case, the SEV is 1, which means that simply changing from having no credit history to having a credit history can make the prediction flip, despite the other factors, and regardless of the total number of features used to construct the model. If most of the feature values are 0 for a given individual (perhaps most people do not have complicated financial histories), or if most factors are not crucial for the prediction, the global sparsity of the model is essentially irrelevant. Hence, SEV is an alternative way of measuring sparsity for these cases.
This work introduces the SEV and discusses some of its properties. SEV (as well as global sparsity) are most relevant for tabular datasets where each feature is meaningful. We show that decision sparsity, as measured by SEV, is naturally low in most common types of machine learning models already, without modification. This means that despite the complexity of these models, their decisions are often based only on a few factors. While SEV is useful as a post-hoc explanation method, it begs the question of whether models with better prediction-level sparsity can be created. In other words, if we care about sparse explanations, we should target them directly instead of obtaining globally sparse models and hoping they give the sparsest explanations. Our second contribution is to create models that are optimized to yield sparse (i.e., low-SEV) explanations. We do this by introducing SEV loss terms that can be used with a variety of models (linear, boosted trees, multi-layer perceptrons, etc.). SEV can also be used with restrictions on all those models to accommodate features that cannot be used as part of the decision-sparsity calculation. The application of our SEV optimization methods to a variety of real-world datasets in our experiments demonstrates that they are extremely effective in improving decision sparsity.
@@ 2 Related Works
Global Model Sparsity. Loosely, globally sparse models (e.g., Ravikumar et al., 2009; Wen et al., 2016; Filzmoser et al., 2012; Tibshirani, 1996; Bach et al., 2012; Zhang et al., 2015) do not have many “components,” such as non-zero coefficients for additive models or leaves of a decision tree. Our focus in this paper is not the overall complexity of a model, but rather the complexity of the individual predictions it generates. These two notions of sparsity — global and local — are undoubtedly related. Even if every variable in a sparse model is used in every prediction, each explanation is necessarily sparse. However, there non-sparse models can yield lower SEVs, as shown in Section 5.1.
@@
There has been a large push to develop methods to explain the predictions of black boxes (e.g., Guidotti et al., 2018; Ribeiro et al., 2016, 2018; Lundberg and Lee, 2017; Baehrens et al., 2010). Those methods evaluate the contribution of features to the final outcome of the model. Some works are limited to particular domains and data types, such as images (e.g., Apicella et al., 2019, 2020; Boreiko et al., 2022), or text (e.g., Lei et al., 2016; Li et al., 2016; Treviso and Martins, 2020; Bastings et al., 2019; Yu et al., 2019, 2021). Chen et al. (2018) provides a subset features selection algorithm called Learning to Explain (L2X), but with an entirely different premise – that sparsity values are fixed by the user and not adaptive. A more detailed comparison can be found in Appendix H.
@@
Counterfactual explanations and algorithmic recourse study what could have happened if input to a model had been changed. Some methods are limited to text (Martens and Provost, 2014), or specific models, like linear models (Ustun et al., 2019), tree ensembles (Cui et al., 2015; Tolomei et al., 2017; Lucic et al., 2022), support vector machines (Gupta et al., 2019), or neural networks (Zhang et al., 2018). Most research in counterfactual explanations or algorithmic recourse focuses on minimizing $\ell_{2}$ distance (Ross et al., 2021), $\ell_{1}$ distance (Wachter et al., 2017; Zhang et al., 2018; Mc Grath et al., 2018; Russell, 2019; Joshi et al., 2019), cosine distance (Ramon et al., 2020), the overall cost of changing features by various amounts, a combination of distances (Laugel et al., 2018; Dhurandhar et al., 2018; Van Looveren and Klaise, 2021; Cheng et al., 2020) or model function outputs (Lash et al., 2017a, b) between the instance’s x value and the closest point across the decision boundary. An important distinction is that we use $\ell_{0}$ distance between explanations to measure sparsity. The only paper we know of that uses $\ell_{0}$ distance is that of Fernández-Loría et al. (2022), which is a special case of our ${}^{-}$; Fernández-Loría et al. (2022) does not discuss sparsity levels for the overall model. As discussed, prediction sparsity is important: an explanation for recourse might involve many changes to someone’s credit and lifestyle since the shortest path to crossing the decision boundary generally involves changes to many variables. Our explanations are sparse (e.g., “you have no credit history”), so they are easier to understand.
@@
Additionally, there is ample evidence from human studies that the counterfactual explanations generated by most methods are not interpretable to humans because they aim only to cross the decision boundary, where the decision made by the model is not obvious to a human. Extensive experiments of Delaney et al. (2023) indicate that minimal changes to an observation that leads to a change in the predicted class are not natural for humans because the decision boundary is crossed in a low-density part of the space, where the observation is not clearly one class or the other. By framing SEV with respect to a population attribute reference (i.e., typical values of features), it provides explanations that humans are more likely to understand. Importantly, Delaney et al. (2023) show that counterfactual explanations created by humans are not close to the decision boundary; people create counterfactuals that appear to be prototypes of the counterfactual class. Our best competitor (counterfactual explanation method DiCE) has precisely this problem, shown in Appendix E.
@@ 3 Sparse Explanation Values
Let us define SEV. We are in the setting of classification, where we have units with features $\boldsymbol{x}\in\mathcal{X}\subseteq\mathbb{R}^{p}$ and an outcome of interest. Here, we consider binary outcomes $\{0,1\}$, though the extension to multi-class outcomes is straightforward. We predict the outcome using a classifier $f:\mathcal{X}\rightarrow\{0,1\}$ that maps the input $\boldsymbol{x}$ to a binary prediction $\hat{y}$. We consider explanations for only one class; to consider explanations for the other class, one would use a completely symmetric procedure (defining a reference for the other class). Without loss of generality, we assume class 1 is the class of interest (positive) and we focus on explaining the predictions of a unit (which we call a query, denoted $\boldsymbol{x}^{\textrm{query}}$) that is classified as positive ($f(\boldsymbol{x}^{\textrm{query}})=1$). As shown by human studies (Delaney et al., 2023), it is natural to compare a query with a reference; for example, a unit representative of the population as a whole or of negatively predicted (predicted to be class 0) units. Therefore, we also define explanation sparsity with respect to the reference $\boldsymbol{\tilde{x}}\in\mathbb{R}^{p}$, where $\tilde{x}_{j}$ is the reference value of feature $j$, e.g., the population mean or median of that feature in the dataset. We will consider aligning the observed values of the query $\boldsymbol{x}_{j}^{\textrm{query}}$ with the corresponding reference feature values $\tilde{\boldsymbol{x}}_{j}$ to see whether predictions change. Thinking of these adjustments as binary moves, it is convenient to represent the $2^{p}$ possible different feature sets as vertices lying on a Boolean hypercube. The hypercube is defined below:
@@
The SEV hypercube definition can also be extended from a hypercube to a Boolean lattice as they have the same geometric structure. Based on this definition of SEV hypercubes, we define two variants of the Sparse Explanation Value: one is defined as the smallest number of features that, starting from the reference (corresponding to the origin of the hypercube $\boldsymbol{b}=\mathbf{0}$), need to be aligned with the query’s feature values to make the prediction flip from negative to positive (called ${}^{+}$). The other (called ${}^{-}$) is the counterfactual case: it is the minimal number of features, starting from the query (corresponding to $\boldsymbol{b}=\mathbf{1}$) that need to be aligned with the reference’s feature values to make the prediction flip from positive to negative. Formally:
@@
${}^{-}$ is similar to a counterfactual explanation in which we use the length of the path on the SEV hypercube to measure the counterfactual change (see also Section 5.4). Figure 0(a), 0(b) and Table 0(c), 0(d) show examples of how ${}^{+}$$=1$ and ${}^{-}$$=2$ are calculated in a loan credit evaluation decision involving three features: housing = renting, loan amount = >10k, and education level = high school level, where we need an SEV hypercube with $82^{3}=$ 8 vertices. The red vertex in Figure 0(a) is encoded as $(1,1,1)$, which represents the query; the dark blue vertex is encoded as $(0,0,0)$, representing the negative-predicted reference value; the orange vertices are positively-predicted vertices; and the light blue ones are negatively predicted vertices. To compute ${}^{+}$, we start from the reference $(0,0,0)$ and find the shortest path to any positive (red/orange) nodes. Here, the two shortest paths end at $(0,1,0)$, which means that if we change the loan amount from 5-10k to more than 10k, then the model will evaluate it as high risk. Since we are changing only one feature (loan amount), ${}^{+}$ equals 1, and the feature vector that $(0,1,0)$ represents is called an ${}^{+}$ explanation. For ${}^{-}$, we start from the query $(1,1,1)$, and find the shortest path to any blue node. Here the shortest paths have length 2 (e.g., $(1,1,1)\to(0,1,1)\to(0,0,1)$), which means that if we changed the housing status of the renter from renting to managing and the loan amount from greater than 10k to 5-10k, then the credit risk will be evaluated as low instead of high. Moreover, there also exists other shortest paths like $(1,1,1)\to(1,1,0)\to(1,0,0)$ which means we can also align the loan amount and the education level with their reference values to obtain a low-risk prediction. This alternative path to a negative vertex shows that we are able to provide multiple sparse explanations through the SEV hypercube. The feature vector that $(1,0,0)$ represents is called an ${}^{-}$ explanation.
Figures 0(e) and 0(f) show how SEV is distributed in 2D cases. The blue line is the decision boundary. The orange horizontal and vertical lines are where the decision boundary intersects one of the reference’s two feature values. If the query’s feature values are in the yellow area of the ${}^{+}$ figure (Figure 0(e)), it means that for both features, the reference feature value must be changed to the query’s feature values in order to be predicted positive. The yellow region in Figure 0(f) (${}^{-}$ case) shows that for both features, the query’s feature values must be align with the reference values in order to be predicted as negative.
@@
${}^{+}$, ${}^{-}$, and ${}^{\circledR}$ can all be computed by a breadth-first search along the SEV hypercube. Since we care about sparse and faithful explanations (i.e., no more than 5-7 features) we can set up search depth limits, which also have the effect of lowering the computational time. SEVs are usually very low in practice, and thus can generally be computed in milliseconds. Detailed results are shown in Section 5.1. Appendix D also shows that SEV requires equal or less time to generate explanations compared to other methods.
@@ 4 Optimizing Sparse Explanation Values
In this section, we describe how we are able to optimize the average ${}^{-}$ and ${}^{+}$ over data points yielded by our classifiers. As we mentioned, calculating SEVs is theoretically a combinatorial problem, and thus directly optimizing it is intractable. However, we propose some easy-to-optimize surrogate objectives that obtain classifiers with very low SEVs in practice.
4.1 Volume-based SEV Plus Optimization for Linear Models

### 09 Mar 2024 v3

@@ Abstract
Even if a model is not globally sparse, it is possible for decisions made from that model to be accurately and faithfully described by a small number of features. For instance, an application for a large loan might be denied to someone because they have no credit history, which overwhelms any evidence towards their creditworthiness. In this work, we introduce the Sparse Explanation Value (SEV), a new way of measuring sparsity in machine learning models. In the loan denial example above, the SEV is 1 because only one factor is needed to explain why the loan was denied. SEV is a measure of decision sparsity rather than overall model sparsity, and we are able to show that many machine learning models – even if they are not sparse – actually have low decision sparsity, as measured by SEV. SEV is defined using movements over a hypercube, allowing SEV to be defined consistently over various model classes, with movement restrictions reflecting real-world constraints. We propose algorithms that reduce SEV without sacrificing accuracy, providing sparse and completely faithful explanations, even without globally sparse models.
@@ 1 Introduction
As machine learning is increasingly leveraged in business and societal contexts to make important decisions, there is an increasing need for insight into why these decisions were made. One classical important measure of interpretability is sparsity; famously, humans can handle only $7\pm 2$ cognitive entities at once (Miller, 1956). Traditionally, sparsity is a property of the entire model, which we call global sparsity, for instance, the total number of terms in linear models, the total number of parameters in neural networks, or the number of leaves in decision trees (Murdoch et al., 2019). In this work, we argue that this notion of sparsity, while undeniably useful, may be overly restrictive. This is because we do not need the model to be globally sparse in order to generate a sparse explanation for use in decision-making. A far more relevant desideratum for users is that the explanation for each individual prediction is sparse. By this, we mean that the local explanation for a specific prediction depends only on a small number of features, which may differ across units. Since individuals care only about decisions made about themselves, we should consider aiming for sparse explanations – and not necessarily sparse models. That is, having an explanation for each prediction that does not necessarily contain all factors used by the model. Having sparse explanations, where we might faithfully explain a given prediction using only 1-3 features, allows for increased transparency in model-assisted decision-making processes, even if the model is not globally sparse.
@@
For example, consider a scenario where a loan applicant is denied a large loan due to their lack of credit history. This denial might occur even though the applicant has several positive factors, such as having a bank account, being employed, and not having a criminal record. In this case, the SEV is 1, which means that simply changing from having no credit history to having a credit history can flip the prediction, despite the other factors, and regardless of the total number of features in the model. If most of the feature values are 0 for a given individual (perhaps most people do not have complicated financial histories), or if most predictions depend only on a small set of factors (where important factors can differ across predictions), the global sparsity of the model is essentially irrelevant. Hence, SEV is an alternative way of measuring sparsity for these cases.
This work introduces the SEV and discusses its properties. SEV (as well as global sparsity) are most relevant for tabular datasets where each feature is meaningful. We show that decision sparsity, as measured by SEV, is naturally low in most common types of machine learning models already, without modification. This means that despite the complexity of these models, their decisions are often based only on a few factors. While SEV is useful as a post-hoc explanation method, it begs the question of whether models with better prediction-level sparsity can be created. In other words, if we care about sparse explanations, we should target them directly instead of obtaining globally sparse models and hoping they give the sparsest explanations. Our second contribution is to create models that are optimized to yield sparse (i.e., low-SEV) explanations. We do this by introducing SEV loss terms that can be used with a variety of models (linear, boosted trees, multi-layer perceptrons, etc.). Using real-world datasets, we show that SEV optimization is effective in improving decision sparsity.
@@ 2 Related Works
Global Model Sparsity. Loosely, globally sparse models (e.g., Rudin et al., 2022; Filzmoser et al., 2012; Tibshirani, 1996; Bach et al., 2012; Zhang et al., 2015) do not have many “components,” such as non-zero coefficients for additive models or leaves of a decision tree. Our focus in this paper is not overall model complexity, but rather the complexity of the individual predictions. These two notions of sparsity — global and local — are undoubtedly related. Even if every variable in a sparse model is used in every prediction, each explanation is necessarily sparse. However, non-sparse models can yield lower SEVs, as shown in Section 5.1.
@@
There has been a large push to develop methods to explain the predictions of black boxes (e.g., Guidotti et al., 2018; Ribeiro et al., 2016, 2018; Lundberg and Lee, 2017; Baehrens et al., 2010). Those methods evaluate the contribution of features to the final outcome of the model. Some works are limited to particular domains and data types, such as images (e.g., Apicella et al., 2019, 2020; Boreiko et al., 2022), or text (e.g., Lei et al., 2016; Li et al., 2016; Treviso and Martins, 2020; Bastings et al., 2019; Yu et al., 2019, 2021). Chen et al. (2018) provides a subset features selection algorithm called Learning to Explain (L2X), but with an entirely different premise – that sparsity values are fixed by the user and not adaptive. A more detailed comparison can be found in Appendix G.
@@
Counterfactual explanations and algorithmic recourse study what could have happened if input to a model had been changed. Some methods are limited to text (Martens and Provost, 2014), or specific model classes, like linear models (Ustun et al., 2019), tree ensembles (Cui et al., 2015; Tolomei et al., 2017; Lucic et al., 2022), support vector machines (Gupta et al., 2019), or neural networks (Zhang et al., 2018). Most research in counterfactual explanations or algorithmic recourse focuses on minimizing $\ell_{2}$ distance (Ross et al., 2021), $\ell_{1}$ distance (Wachter et al., 2017; Zhang et al., 2018; McGrath et al., 2018; Russell, 2019; Joshi et al., 2019), cosine distance (Ramon et al., 2020), the overall cost of changing features by various amounts, a combination of distances (Laugel et al., 2018; Dhurandhar et al., 2018; Van Looveren and Klaise, 2021; Cheng et al., 2020) or model function outputs (Lash et al., 2017a, b) between the instance’s x value and the closest point across the decision boundary. An important distinction is that we use $\ell_{0}$ distance between explanations to measure sparsity. The only paper we know of that uses $\ell_{0}$ distance is that of Fernández-Loría et al. (2022), which is a special case of our ${}^{-}$; Fernández-Loría et al. (2022) does not discuss sparsity as an evaluation metric for explanations. As discussed, prediction sparsity is important: a classic counterfactual explanation might involve many changes to someone’s credit and lifestyle since the shortest path to crossing the decision boundary generally involves changes to many variables. Our explanations are sparse (e.g., “you have no credit history”), so they are easier to understand.
@@
Additionally, there is ample evidence from human studies that the counterfactual explanations generated by most methods are not interpretable to humans because they aim only to cross the decision boundary, where the decision made by the model is not obvious to a human. Extensive experiments of Delaney et al. (2023) indicate that minimal changes to an observation that alter the predicted class are not natural for humans because the decision boundary is crossed in a low-density part of the space where the observation is not clearly one class or the other. Our best competitor (counterfactual explanation method DiCE) has precisely this problem, shown in Section 5.4. By framing SEV with respect to a population attribute reference (i.e., typical values of features), it provides explanations that humans are more likely to understand. Importantly, Delaney et al. (2023) show that counterfactual explanations created by humans are not close to the decision boundary; people create counterfactuals that appear to be class-specific prototypes.
@@ 3 Sparse Explanation Values
Let us define SEV. We are in the setting of classification, where we have observations (units) with features $\boldsymbol{x}\in\mathcal{X}\subseteq\mathbb{R}^{p}$ and an outcome of interest. Here, we consider binary outcomes $\{0,1\}$, though the extension to multi-class outcomes is straightforward. We predict the outcome using a classifier $f:\mathcal{X}\rightarrow\{0,1\}$ that maps the input $\boldsymbol{x}$ to a binary prediction $\hat{y}$. We consider explanations for only one class; to consider explanations for the other class, one would use a completely symmetric procedure (defining a reference for the other class). Without loss of generality, we assume class 1 is the class of interest (positive) and we focus on explaining the predictions of a unit (which we call a query, denoted $\boldsymbol{x}^{\textrm{query}}$) that is classified as positive ($f(\boldsymbol{x}^{\textrm{query}})=1$). As shown by human studies (Delaney et al., 2023), it is natural to compare a query with a reference; for example, a unit representative of the population as a whole or of negatively predicted (predicted to be class 0) units. Therefore, we define explanation sparsity with respect to the reference $\boldsymbol{\tilde{x}}\in\mathbb{R}^{p}$, where $\tilde{x}_{j}$ is the reference value of feature $j$, e.g., the population mean or median of that feature in the dataset. We will consider aligning the observed values of the query $\boldsymbol{x}_{j}^{\textrm{query}}$ with the corresponding reference feature values $\tilde{\boldsymbol{x}}_{j}$ to see whether predictions change. Thinking of these adjustments as binary moves, it is convenient to represent all $2^{p}$ possible different feature sets as vertices of on a Boolean hypercube. The hypercube is defined below:
@@
The SEV hypercube definition can also be extended from a hypercube to a Boolean lattice as they have the same geometric structure. Based on this definition of SEV hypercubes, we define two variants of the Sparse Explanation Value: one is the smallest number of features that, starting from the reference (corresponding to the origin of the hypercube $\boldsymbol{b}=\mathbf{0}$), need to be aligned with the query’s feature values to make the prediction flip from negative to positive (called ${}^{+}$). The other (called ${}^{-}$) is the counterfactual case: it is the minimal number of features, starting from the query (corresponding to $\boldsymbol{b}=\mathbf{1}$) that need to be aligned with the reference’s feature values to make the prediction flip from positive to negative. Formally:
@@
${}^{-}$ is similar to a counterfactual explanation in which we use the length of the path on the SEV hypercube to measure the counterfactual change (see also Section 5.4). Figure 0(a), 0(b) and Table 0(c), 0(d) show examples of how ${}^{+}$$=1$ and ${}^{-}$$=2$ are calculated in a loan credit evaluation decision involving three features: housing = renting, loan amount $\geq$ 10k, and education level = high school level, where we need an SEV hypercube with $82^{3}=$ 8 vertices. The red vertex in Figure 0(a) is encoded as $(1,1,1)$, which represents the query; the dark blue vertex is encoded as $(0,0,0)$, representing the negatively-predicted reference value; the orange vertices are positively-predicted vertices; and the light blue ones are negatively predicted vertices. To compute ${}^{+}$, we start from the reference $(0,0,0)$ and find the shortest path to any positive (red/orange) nodes. Here, the two shortest paths end at $(0,1,0)$, which means that if we change the loan amount from 5-10k to more than 10k, then the model will evaluate it as high risk. Since we are changing only one feature (loan amount), ${}^{+}$ equals 1, and the feature vector that $(0,1,0)$ represents is called an ${}^{+}$ explanation. For ${}^{-}$, we start from the query $(1,1,1)$, and find the shortest path to any blue node. Here the shortest paths have length 2 (e.g., $(1,1,1)\to(0,1,1)\to(0,0,1)$), which means that if we changed the housing status of the renter from renting to managing and the loan amount from greater than 10k to 5-10k, then the credit risk will be evaluated as low instead of high. Moreover, there also exists other shortest paths like $(1,1,1)\to(1,1,0)\to(1,0,0)$ which means we can also align the loan amount and the education level with their reference values to obtain a low-risk prediction. This alternative path to a negative vertex shows that we are able to provide multiple sparse explanations through the SEV hypercube. The feature vector that $(1,0,0)$ represents is called an ${}^{-}$ explanation.
Figures 0(e) and 0(f) show how SEV is distributed in 2D cases. The blue line is the decision boundary. The orange horizontal and vertical lines are where the decision boundary intersects one of the reference’s two feature values. If the query’s feature values are in the yellow area of the ${}^{+}$ figure (Figure 0(e)), it means that for both features, the reference feature value must be changed to the query’s feature values in order to be predicted positive. In the yellow region of Figure 0(f) (${}^{-}$ case), for both features, the query’s feature values must be aligned with the reference values in order to be predicted as negative.
@@
${}^{+}$, ${}^{-}$, and ${}^{\circledR}$ can all be computed by a breadth-first search along the SEV hypercube. Since we care about sparse and faithful explanations (i.e., no more than 5-7 features) we can set up search depth limits, which also have the effect of lowering the computational time. SEVs are usually very low in practice, and thus can generally be computed in milliseconds. Detailed results are shown in Section 5.1. Appendix D also shows that SEV requires equal or less time to generate explanations than other methods.
@@ 4 Optimizing Sparse Explanation Values
In this section, we describe how to obtain classifiers that optimize the average ${}^{-}$ or ${}^{+}$. As we mentioned, calculating SEVs is theoretically a combinatorial problem. Thus directly optimizing it is intractable. Therefore, we propose some easy-to-optimize surrogate objectives that obtain classifiers with very low SEVs in practice.
4.1 Volume-based ${}^{+}$ Optimization for Linear Models